OpenStudy (yacoub1993):

You have a coupon good for $6 off the price of any large pizza. You also get a 20% discount on any pizza if you show your student ID. How much more would you pay for a large pizza if the cashier applies the coupon first? A. $1.20 B. $0.00 C. $1.50 D. $.50

OpenStudy (yacoub1993):

D

OpenStudy (yacoub1993):

@BangkokGarrett

OpenStudy (anonymous):

Say a pizza is normally $10 $6 off first: now its $4 20% discount: 4 - 0.2*4 = 3.20 ---- 20% off first: 10-0.2*10 = $8 $6 off: $2 --- 3.2 - 2 = 1.20